OpenStudy (anonymous):

Steps on how to establish the identity: (cos²θ/cosθ-sinθ) + (sin²θ/sinθ-cosθ) = sinθ+cosθ How can I simplify the left side even further?

zepdrix (zepdrix):

\[\large \frac{\cos^2\theta}{\cos \theta-\sin \theta}+\frac{\sin^2\theta}{\sin \theta-\cos \theta}=\sin \theta+\cos \theta\] Let's start by factoring out a negative -1 from the right fraction. This might not be immediately obvious why I did this. Let me explain: We want to be able to add these fractions together. To do so, we need to have a common denominator.\[\large \frac{\cos^2\theta}{\cos \theta-\sin \theta}+\frac{\sin^2\theta}{\sin \theta-\cos \theta} \qquad \rightarrow \qquad \frac{\cos^2\theta}{\cos \theta-\sin \theta}+\frac{\sin^2\theta}{-(\cos \theta-\sin \theta)}\] \[\large \rightarrow \qquad \frac{\cos^2\theta}{\cos \theta-\sin \theta}-\frac{\sin^2\theta}{\cos \theta-\sin \theta}\] Understand what I did there? That step can be a little tricky.

OpenStudy (anonymous):

@zepdrix Thank you so much, I got it! My textbook completely skipped this step in the answer key, so I didn't really know how to proceed at that point. Thank you so much! :)

zepdrix (zepdrix):

Oh cool! c: